A bond buyer paid 900 for a bond that will be valued at 1000$ at maturity, in three months. What is the annualized rate of return on this bond?
Answer #1

Buy Price of Bond = $900

In 3 months, bond will mature at $1000

Annualized Rate of return = (\frac{Maturity Value}{Buy Price})^{\frac{1}{n}}-1

where, n = no of periods = 3 months/12 months = 0.25

Annualized Rate of return = (\frac{1000}{900})^{\frac{1}{0.25}}-1

= 1.524158 - 1

= 52.4158%

So, the annualized rate of return on this bond is 52.4158%